# Towards Task-Compatible Compressible Representations

### v2

@@ 3.2 Task-Compatible Representations
$\sup_{\mathbf{f}\in\mathcal{F}}\left\{I_{\mathcal{V}_{b}}(\mathbf{f}(X)\to Z_{b})+\beta I_{\mathcal{V}_{a}}(\mathbf{f}(X)\to X)\right\},$ (2)
where $\beta$ is a sufficiently small positive scalar and $\mathcal{V}_{a}$ is a predictive set given by the set of synthesis transforms $\mathcal{G}_{a}$. By adding a small reward on the capacity of $\mathbf{f}$ to reconstruct the input using a simple transform, we prevent the co-adaptation or specialization of $Y$ to the main task $Z_{b}$. The choice of reconstruction as a task and a simple synthesis transform ensures that an approximation of the original input can be recovered by other tasks. This seems to be a good general choice, since task models are often designed to work on the original input space and can be trained in an unsupervised manner. We design the architecture defining $\mathcal{G}_{a}$ to be subsumed by the synthesis transforms $\mathcal{G}_{e}$ of the secondary task, ensuring that the model has enough expressive power to reconstruct the input and perform the task, if necessary.
@@ 3.5 Scalable Compression
$\mathcal{L}_{b}(\psi_{b},\phi_{b},\theta_{a})=\lambda_{b}\mathbb{E}[d_{b}(\hat{Z}_{b},Z_{b})]+R_{b}+\beta\,\mathbb{E}[d_{a}(\hat{Z}_{a},X)];$
$Y_{b}=\mathbf{f}_{b}(X;\psi_{b}),\hat{Z}_{b}=\mathbf{g}_{b}(\hat{Y}_{b};\theta_{b}),\hat{Z}_{a}=\mathbf{g}_{a}(\hat{Y}_{b};\theta_{a}),$
@@
where $\hat{h}(\cdot;\phi_{b})$ is an entropy estimate parameterized by $\phi_{b}$, $d_{b}(\cdot,\cdot)$ is the base task loss, $d_{a}(\cdot,\cdot)$ is the reconstruction distortion, and $\lambda_{b}$ and $\beta$ are hyper-parameters. The $\lambda$ hyper-parameter controls the rate-distortion trade-off. For the enhancement task, we use the traditional rate-distortion loss function:
@@
where $\hat{h}(\cdot|\cdot;\phi_{e})$ is the conditional entropy estimate, parameterized by $\phi_{e}$. During training, either $\mathbf{f}_{e}(\cdot;\psi_{b})$ remains frozen, or the gradients from the enhancement loss do not flow to it.
@@ 4 Experiments
Fig. 2: Rate-distortion performance of base representations on image reconstruction (a)-(d) and base tasks (e)-(f). Left: object detection on COCO 2017; Right: depth estimation on Cityscapes. “Direct” means reconstructing the input from the base representation directly, without the dedicated channel (see Fig. 1). “Scalable” means conditionally coding reconstruction representation based on the base representation; in this case, the BPP reported is the sum of both representations. The “Standalone” method does not use side-information. The lines marked “Uncompressed” correspond to the best task performance obtained with a very large $\lambda_{b}=1,000$, amongst the baseline and proposed methods. PSNR units are in decibels (dB).
@@ 5 Conclusion
The $\mathcal{V}$-information framework provides a possible interpretation in terms of the increase of predictive $\mathcal{V}$-information in the base representation $Y_{b}$ for a predictive family $\mathcal{V}_{e}$. This would especially be the case when the auxiliary task for $Y_{b}$ is sufficiently non-specific (such as input reconstruction), and $\mathcal{V}_{a}$ is sufficiently simple, so as to reduce co-adaptation and be able to generalize to other tasks.

### v3

@@ 3.2 Task-Compatible Representations
$\sup_{\mathbf{f}\in\mathcal{F}}\left\{I_{\mathcal{V}_{b}}(\mathbf{f}(X)\to Z_{b})+\beta I_{\mathcal{V}_{a}}(\mathbf{f}(X)\to Z_{a})\right\},$ (2)
where $Z_{a}$ is an auxiliary task, $\beta$ is a sufficiently small positive scalar, and $\mathcal{V}_{a}$ is a predictive set given by the set of synthesis transforms $\mathcal{G}_{a}$. By setting $Z_{a}=X$, we add a small reward on the capacity of $\mathbf{f}$ to reconstruct the input using a simple transform, preventing the co-adaptation or specialization of $Y$ to the main task $Z_{b}$. The choice of reconstruction as a task and a simple synthesis transform ensures that an approximation of the original input can be recovered by other tasks. This seems to be a good general choice, since task models are often designed to work on the original input space and can be trained in an unsupervised manner. We design the architecture defining $\mathcal{G}_{a}$ to be subsumed by the synthesis transforms $\mathcal{G}_{e}$ of the secondary task, ensuring that the model has enough expressive power to reconstruct the input and perform the task, if necessary.
@@ 3.5 Scalable Compression
$\mathcal{L}_{b}(\psi_{b},\phi_{b},\theta_{b},\theta_{a})=\lambda_{b}\mathbb{E}[d_{b}(\hat{Z}_{b},Z_{b})]+R_{b}+\beta\mathbb{E}[\|{\hat{X}\shortminus X\|}];$
$Y_{b}=\mathbf{f}_{b}(X;\psi_{b}),\hat{Z}_{b}=\mathbf{g}_{b}(\hat{Y}_{b};\theta_{b}),\hat{X}=\mathbf{g}_{a}(\hat{Y}_{b};\theta_{a}),$
@@
where $\hat{h}(\cdot;\phi_{b})$ is an entropy estimate parameterized by $\phi_{b}$, $d_{b}(\cdot,\cdot)$ is the base task loss. $\lambda_{b}$ and $\beta$ are hyper-parameters. The $\lambda$ hyper-parameter controls the rate-distortion trade-off. For the enhancement task, we use the traditional rate-distortion loss function:
@@
where $\hat{h}(\cdot|\cdot;\phi_{e})$ is the conditional entropy estimate, parameterized by $\phi_{e}$. During training, either $\mathbf{f}_{b}(\cdot;\psi_{b})$ remains frozen, or the gradients from the enhancement loss do not flow to it.
@@ 4 Experiments
Fig. 2: Rate-distortion performance of base representations on image reconstruction (a)-(d) and base tasks (e)-(f). Left: object detection on COCO 2017; Right: depth estimation on Cityscapes. “Direct” means reconstructing the input from the base representation directly, without the dedicated channel (see Fig. 1). “Scalable” means coding reconstruction representation conditional on the base representation; in this case, the BPP reported is the sum of both representations. The “Standalone” method does not use side-information. The lines marked “Uncompressed” correspond to the best task performance obtained with a very large $\lambda_{b}=1,000$, amongst the baseline and proposed methods. PSNR units are in decibels (dB).
@@ 5 Conclusion
The $\mathcal{V}$-information framework provides a possible interpretation in terms of the increase of predictive $\mathcal{V}$-information in the base representation $Y_{b}$ for a predictive family $\mathcal{V}_{e}$. This would especially be the case when the auxiliary task inducing $Y_{b}$ is sufficiently non-specific (such as input reconstruction), and $\mathcal{V}_{a}$ is sufficiently simple, so as to reduce co-adaptation and generalize to other tasks.